A lady is weighing some bananas in a grocery store when the floor

collapses. If the banana's mass is 2 kg and the floor is accelerating

at -2.25 m/s2, what is the apparent weight (normal force) of the

bananas?


1
Expert's answer
2021-11-11T17:07:56-0500

The apparent weight is


"W_a=m(g+a)=2(9.8-2.25)=15.1\text{ N},"

which corresponds to apparent mass of


"m_a=W_a/g=1.54\text{ kg}."